OpenStudy (anonymous):

Match the equation on the right; 1 + sec^2 x sin^2 x = sec^2 x

OpenStudy (anonymous):

I don't need the answer, I just need someone to start me off, please.

OpenStudy (anonymous):

First you would divide Sec^2 from the left side. You would be left with 1 + X sin^2 X = X

OpenStudy (anonymous):

X sin^2 X = X - 1 Then simplify it to 2x sin^2 = X - 1 Then get rid of the 2x and simplify it some more..